OpenStudy (anonymous):

The displacement from equilibrium of an object in harmonic motion on the end of a spring is: y= (1/3)cos12t-(1/4)sin12t where y is measured in feet and t is the time in seconds. Determine the position and velocity of the object when t= (pi/8) ?

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@kelliegirl33 @RBauer4

OpenStudy (anonymous):

i got this far f(x) = g(x)h(x) f(x)' = g(x)' (h(x)) + h(x)' (g(x)) f(x) = (1/3)cos12t-(1/4)sin12t g(x) = (1/3)cos12t-(1/4) -- You have to apply chain rule to this function when finding the derivative h(x) = sin12t f(x)' = -(1/3)sin(12t - (1/4) * ( 12) *(sin12t) + (cos12t)(12)*((1/3)cos12t-(1/4))